# Private Online Community Detection for Censored Block Models

### v1

@@ Abstract
We study the private online change detection problem for dynamic communities, using a censored block model (CBM). Focusing on the notion of edge differential privacy (DP), we seek to understand the fundamental tradeoffs between the privacy budget, detection delay, and exact community recovery of community labels. We establish the theoretical lower bound on the delay in detecting changes privately and propose an algorithm capable of identifying changes in the community structure, while maintaining user privacy. Further, we provide theoretical guarantees for the effectiveness of our proposed method by showing necessary and sufficient conditions on change detection and exact recovery under edge DP. Simulation and real data examples are provided to validate the proposed method.
1 Introduction
@@
Community detection in networks is a fundamental problem in the area of graph mining and unsupervised learning [14]. The main goal is to find partitions (or communities) in the graph that have dense connections within communities and sparse connections across communities. The Stochastic Block Model (SBM) is commonly used to model the graphical structure for networks [2, 17]. In SBM model, nodes are divided into communities, and the connection probability between nodes depends on whether they belong to the same communities or not. Connections between nodes are more probable within the same community than between communities. This difference in connection probabilities is central to the challenge of community detection. The study and improvement of community detection methods using the SBM framework has been a dynamic area of research, with many advances and findings detailed in comprehensive surveys such as the one by Abbe et al. [1].
@@
In contemporary research on network community detection, a significant focus has been placed on static network structures. However, this focus on static networks does not completely capture the dynamic nature of real-world networks, where the arrangement of nodes and their links constantly changes. Despite some advances (e.g., [37, 3]) in applying the dynamic SBM to these scenarios, a comprehensive understanding of the conditions that allow for accurate community detection in such dynamic environments, particularly within the framework of the dynamic Censored Block Model (CBM), remains an open and critical area of research.
@@
The study of CBM in dynamical networks has become increasingly relevant in network science, offering critical insights into the dynamics of complex systems [16, 25]. CBM models effectively represent real-world networks characterized by block structures and are particularly valuable in analyzing networks with incomplete or censored data. This feature is essential for understanding various network dynamics, such as information spread and epidemic outbreaks. CBM models are also instrumental in the development and testing of network science algorithms, especially for community detection and network reconstruction.
@@
The information-theoretic conditions for recoverability are well understood in terms of the scaling of the CBM parameters $p$ and $\zeta\in(0,1/2)$, where $p$ represents the probability of an edge being observed, while $\zeta$ denotes the connection probability (these quantities will be defined more precisely in the sequel). In particular, in the dense regime (which is the focus of this paper), with $p=a\log(n)/n$ for some constant $a>0$, it is known that exact recovery under binary CBMs is possible if and only if $a(\sqrt{1-\zeta}-\sqrt{\zeta})^{2}>1$ [18]. The condition above is considered to be a sharp threshold for recovery, which can also be interpreted as a phase transition. In the literature, efficient algorithms for recovering communities have been developed using spectral methods (e.g., [10, 9]) and semi-definite programming (SDP) [17]. Some of these results were proven to achieve the optimal threshold for exact recovery. In contrast, the sparse regime, defined by $p=a/n$, permits only partial recovery, meaning the estimator is positively correlated with the ground truth labels, provided that $a(1-2\zeta)^{2}>1$, as demonstrated in [21]. Furthermore, spectral algorithms have been shown to achieve this sharp threshold efficiently, as evidenced in [31].
@@
In this work, we study the dynamic CBM models under privacy constraints, since sensitive information is frequently present in network data. As such, ensuring the privacy of individuals during data analysis is imperative. Differential Privacy (DP) [12] has become the recognized standard for delivering robust privacy assurances. DP guarantees that any single user’s participation in a dataset only minutely impacts the statistical outcomes of queries. In the literature [20], two main privacy notions have been proposed for data analysis over graphs: $(1)$ edge DP that aims to protect the individual relationships (edges) in a graph, using randomized algorithms, and $(2)$ node DP that focuses of the privacy of the nodes and their corresponding interactions (edges). Notably, DP algorithms have been tailored to specific network analysis problems, including counting stars, triangles, cuts, dense subgraphs, and communities or even generating synthetic graphs, have been introduced [27, 30, 19, 5].
Under edge DP constraints, we study the online change detection of community structures for CBM models. We propose a private online detection algorithm that performs window-based detection on privatized data. Sufficient conditions for exact recovery of the unknown community structure, under DP constraints, are established. Furthermore, we provide a theoretical guarantee for the detection performance of the proposed method and a new fundamental limit to private exact recovery for the community structure of the CBM.
@@
The problem of private change detection is an evolving field with significant applications. It has been studied under $(\epsilon,\delta)$ differential privacy constraints for multiple settings [8, 7]. In [4], the online change detection problem has been studied under the locally private condition. In [22], the authors extended the analysis of privatized networks from static to dynamic, focusing on sequences of networks with change-points. They explored the limits of consistently localizing change points under edge privacy constraints, uncovering a phase transition in the signal-to-noise ratio conditions, and developing polynomial-time algorithms. The research highlighted how privacy influenced the efficiency of change point localization, with a notable difference in the scaling of sparsity parameters compared to non-private scenarios. The algorithms were found to be near-optimal under edge Local Differential Privacy (LDP) constraints. This work underscored the complexities and challenges of preserving privacy in the dynamic analysis of network data.
In contrast, our paper broadens the scope of analysis by considering possible network changes, in particular, shifts in community memberships. This approach presents a more comprehensive and challenging problem, as it involves not only the detection of statistical changes but also the identification of structural and compositional alterations within the network dynamics. Our methodology, therefore, addresses a broader range of changes, offering deeper insights into the complexities of dynamic network analysis.
@@
Boldface uppercase letters denote matrices (e.g., ${\mathbf{A}}$), while boldface lowercase letters are used for vectors (e.g., ${\mathbf{a}}$). We denote scalars by non-boldface lowercase letters (e.g., $x$), and sets by capital calligraphic letters (e.g., ${\mathcal{X}}$). Let $[n]\triangleq[1,2,\cdots,n]$ represent the set of all integers from $1$ to $n$. We use $\operatorname{Bern}(p)$ to denote a Bernoulli random variable with success probability $p$. For asymptotic analysis, we say the function $f(n)=o(g(n))$ when $\lim_{n\rightarrow\infty}f(n)/g(n)=0$. Also, $f(n)=\mathcal{O}(g(n))$ means there exist some constant $C>0$ such that $|f(n)/g(n)|\leq C$, $\forall n$, and $f(n)=\Omega(g(n))$ means there exists some constant $c>0$ such that $|f(n)/g(n)|\geq c$, $\forall n$.
2 Preliminaries
In this section, we first introduce the binary censored block model used to model real-world network structures in Section 2.1. We present useful preliminaries and background on community estimation using maximum likelihood estimator and the corresponding semidefinite relaxation in Section 2.2. The definition of the notion of edge differential privacy used in this work is introduced in Section 2.3.
2.1 Binary Censored Block Model
We consider an undirected graph $G=({V},E)$ consisting of $n$ vertices, where the vertices are divided into two communities $C_{1}$ and $C_{2}$, and ${V}=C_{1}\cup C_{2}$. The community label for vertex $i$ is denoted by $\sigma^{*}_{i}\in\{-1,+1\}$, $\forall i\in[n]$ and we denote the label vector as $\bm{\sigma}^{\ast}=\{\sigma_{1}^{\ast},\ldots,\sigma_{n}^{\ast}\}$. Further, we assume that the graph $G$ is generated through a CBM. The connections between vertices are represented by the adjacency matrix $\mathbf{A}$, where each edge between two vertices $i<j$ is labeled as $A_{i,j}\in\{+1,-1,0\}$ and drawn as follows:
$A_{i,j}\!\sim\!\begin{cases}p(1-\zeta)\delta_{+1}+p\zeta\delta_{-1}+(1-p)\delta_{0},&\text{if }\sigma^{*}_{i}=\sigma^{*}_{j},\\ p(1-\zeta)\delta_{-1}+p\zeta\delta_{+1}+(1-p)\delta_{0},&\text{if }\sigma^{*}_{i}\neq\sigma^{*}_{j},\end{cases}$ (1)
where $p$ represents the revealed probability of an edge. The parameter $\zeta$ lies within the interval $(0,\frac{1}{2})$, and $\delta$ denotes the Dirac delta function, which is centered around its subscript indicating the precise value of a parameter or a variable it is associated with. In our analysis, we consider an undirected graph without self-loops, i.e., $A_{i,i}=0$ for all $i$, ensuring no vertex is connected to itself, and $A_{i,j}=A_{j,i}$ for all $i\neq j$. For simplicity we denote this model as ${\mathsf{CBM}}(\bm{\sigma}^{\ast},p,\zeta)$. In this work, we focus on the dense regime of the CBM, parameterized by $p=a\log(n)/n$ for a positive constant $a$.
The goal of the community detection problem is to design a (stochastic) estimator $\hat{\bm{\sigma}}:\mathbf{A}\rightarrow\{-1,+1\}^{n}$ for community recovery (i.e., estimating the true labeling vector $\bm{\sigma}^{*}$) upon observing the adjacency matrix. We next define the notion of asymptotic exact recovery as a measure of performance of an estimator. Before we introduce the notion, we first present the discrepancy measure between any estimated community labeling vector $\hat{\bm{\sigma}}$ and the ground truth labels $\bm{\sigma}^{*}$ as
$\operatorname{err}(\hat{\bm{\sigma}},{\bm{\sigma}^{*}})$ $=\min\{\operatorname{Ham}(\hat{\bm{\sigma}},{\bm{\sigma}^{*}}),\operatorname{Ham}(-\hat{\bm{\sigma}},{\bm{\sigma}^{*}})\},$ (2)
where $\operatorname{Ham}(\cdot,\cdot)$ denotes the Hamming distance.
Definition 1 (Exact Recovery).
$\operatorname{Pr}(\operatorname{err}(\hat{\bm{\sigma}}(\mathbf{A}),{\bm{\sigma}^{*}})=0)=1-o(1),$ (3)
where the probability is taken over both the randomness of the adjacency matrix ${\mathbf{A}}$ of the observed graph $G$ as well as the stochastic estimation process.
2.2 Maximum Likelihood Estimator and Semidefinite Relaxation
Under the ${\mathsf{CBM}}(\bm{\sigma},p,\zeta)$ model, given an observation ${\mathbf{A}}$, the log-likelihood function can be calculated as (see Appendix A):
$\log p({\mathbf{A}}|\bm{\sigma})=$ $\left(\frac{1}{4}\log\frac{1-\zeta}{\zeta}\right)\bm{\sigma}^{T}{\mathbf{A}}\bm{\sigma}+{n\choose 2}\log(1-p)+E\log\left(\frac{p}{1-p}\sqrt{\zeta(1-\zeta)}\right),$ (4)
where $E=\sum_{i<j}|A_{i,j}|$ denotes the total number of edges for the adjacency matrix ${\mathbf{A}}$.
Therefore, for the binary CBM, given that $\zeta\in(0,\frac{1}{2})$, the Maximum Likelihood Estimator (MLE) of the label vector is the solution to the following optimization problem:
$\hat{\bm{\sigma}}_{\text{ML}}=\arg\max_{\bm{\sigma}}\{\bm{\sigma}^{T}\mathbf{A}\bm{\sigma}:\sigma_{i}=\pm 1\}.$ (5)
Information theoretically, it has been shown in [18] that for CBM within the dense regime, i.e., $p={a\log(n)}/{n}$, exact recovery is possible if and only if $a(\sqrt{1-\zeta}-\sqrt{\zeta})^{2}>1$.
It is worth mentioning that the MLE is equivalent to solving the max-cut problem in the graph, which is known to be NP-hard [18]. Subsequently, several works in the literature have proposed polynomial time algorithms that can achieve the exact recovery. For instance, it has been shown in [18] that semidefinite relaxation of the MLE can also achieve the optimal recovery threshold. We next present the optimization problem of the MLE via semidefinite relaxation as proposed in [18].
Let us first define $\mathbf{Y}={\bm{\sigma}}{\bm{\sigma}}^{T}$, where $Y_{i,i}=1,\forall i\in[n]$. The semidefinite relaxation approach solves the following optimization problem:
$\hat{\mathbf{Y}}=\max_{\mathbf{Y}}$ $\hskip 7.22743pt\text{tr}({\mathbf{A}}\mathbf{Y})$
s.t. $\hskip 7.22743pt\mathbf{Y}\succcurlyeq\mathbf{0},\,Y_{i,i}=1,\forall i\in[n].$ (6)
It has been shown in [18] that the MLE obtained via semidefinite programming achieves exact recovery (after rounding the solution of (6)) if $a(\sqrt{1-\zeta}-\sqrt{\zeta})^{2}>1$.
2.3 Differential Privacy
In addition to community recovery, we require that the recovery algorithm for community detection also protects the individual connections (i.e., the edges in the graph $G$) in the network.
Privacy Threat Model: In the context of community detection within graph data, our threat model is structured around the principle of edge differential privacy (DP), incorporating both local and central settings. Local DP offers robust privacy guarantees during the data collection process, ensuring that individual data contributions are privatized at the source. This is crucial in scenarios where trust in the data curator is limited. Conversely, in the central setting, privacy protections are applied to the aggregated data, focusing on preserving the presence or absence of single edges. The primary sensitivity in this model lies in the relationships between nodes rather than the nodes’ labels. This sensitivity arises from potential privacy risks of exposing how individuals or entities are interconnected, which could reveal undisclosed or sensitive information. The model assumes that an adversary could have partial knowledge about the graph’s structure and seeks to infer additional undisclosed relationships. Our approach aims to ensure that the detection of communities within this graph does not compromise the privacy of the relationships between the nodes in both local and central settings, allowing for meaningful analysis of the overall graph structure while safeguarding individual privacy.
Similar to [22], we consider the privacy mechanisms where each privatized edge $\tilde{\mathbf{A}}_{i,j}$ will depend on the raw edge ${\mathbf{A}}_{i,j}$ through a conditional distribution $Q$, i.e., $\tilde{\mathbf{A}}_{i,j}|{\mathbf{A}}_{i,j}=a\sim Q(\cdot|a)$.

### v2

# Differentially Private Online Community Detection for Censored Block Models: Algorithms and Fundamental Limits

@@ Abstract
We study the private online change detection problem for dynamic communities, using a censored block model (CBM). We consider edge differential privacy (DP) in both local and central settings, and propose joint change detection and community estimation procedures for both scenarios. We seek to understand the fundamental tradeoffs between the privacy budget, detection delay, and exact community recovery of community labels. Further, we provide theoretical guarantees for the effectiveness of our proposed method by showing necessary and sufficient conditions for change detection and exact recovery under edge DP. Simulation and real data examples are provided to validate the proposed methods.
I Introduction
@@
Community detection in networks is a fundamental problem in the area of graph mining and unsupervised learning [17]. The main goal is to find partitions (or communities) in the graph that have dense connections within communities and sparse connections across communities. The Stochastic Block Model (SBM) is commonly used to model the graphical structure for networks [2, 20]. In an SBM model, nodes are divided into communities, and the connection probability between nodes depends on whether they belong to the same community or not. Connections between nodes are more probable within the same community than between communities. This difference in connection probabilities is central to the challenge of community detection. The study and improvement of community detection methods using the SBM framework has been a dynamic area of research, with many advances and findings detailed in comprehensive surveys such as the one by Abbe et al. [1].
@@
In contemporary research on network community detection, a significant focus has been placed on static network structures. However, this focus on static networks does not completely capture the dynamic nature of real-world networks, where the arrangement of nodes and their links may constantly change [41]. Despite some advances (e.g., [3, 40]) in applying the dynamic SBM to these scenarios, a comprehensive understanding of the conditions that allow for accurate community detection in such dynamic environments, particularly within the framework of the dynamic Censored Block Model (CBM), remains an open and critical area of research.
@@
The study of CBMs in dynamical networks has become increasingly relevant in network science, offering critical insights into the dynamics of complex systems [19, 30]. CBM models effectively represent real-world networks characterized by block structures and are particularly valuable in analyzing networks with incomplete or censored data. This feature is essential for understanding various network dynamics, such as information spread and epidemic outbreaks. CBM models are also instrumental in the development and testing of network science algorithms, especially for community detection and network reconstruction.
@@
The information-theoretic conditions for recoverability are well understood in terms of the scaling of the CBM parameters $n$, $p$ and $\zeta\in(0,1/2)$, where $n$ denotes the number of nodes, $p$ represents the probability of an edge being observed, and $\zeta$ denotes the connection probability (these quantities will be defined more precisely in the sequel). In particular, in the dense regime (which is the focus of this paper), with $p=a\log(n)/n$ for some constant $a>0$, it is known that exact recovery under binary CBMs is possible if and only if $a(\sqrt{1-\zeta}-\sqrt{\zeta})^{2}>1$ [21]. This condition is considered to be a sharp threshold for recovery, which can also be interpreted as a phase transition. In the literature, efficient algorithms for recovering communities have been developed using spectral methods (e.g., [11, 12]) and semidefinite programming (SDP) [20]. Some of these results (e.g., [11, 20]) were proven to achieve the optimal threshold for exact recovery.
@@
In this work, we study the dynamic CBM models under privacy constraints, since sensitive information is frequently present in network data. As such, ensuring the privacy of individuals during data analysis is imperative. Differential Privacy (DP) [16] has become the recognized standard for delivering robust privacy assurances. DP guarantees that any single user’s participation in a dataset only minutely impacts the statistical outcomes of queries. In the literature [24], two main privacy notions have been proposed for data analysis over graphs: $(1)$ edge DP that aims to protect the individual relationships (edges) in a graph, and $(2)$ node DP that focuses of the privacy of the nodes and their corresponding interactions (edges). Notably, DP algorithms have been tailored to specific network analysis problems, including counting stars, triangles, cuts, dense subgraphs, and communities, as well as the generation of synthetic graphs [5, 23, 31, 33].
Our privacy threat model incorporates both local and central settings. We highlight that the central DP is a weaker guarantee compared to the local DP (as it can be readily shown that the local DP guarantee implies central DP). The rationale for considering both local and central DP simultaneously is as follows. In scenarios where a trusted curator/server (e.g., Facebook) is initially considered trustworthy, users could send their raw (unperturbed) local connections to the server. However, if the raw connections received by the server were leaked (e.g., via a data breach), then having an additional local DP guarantee can provide a second layer of protection. Furthermore, the local privacy setting has a worse utility guarantee due to the users’ perturbation of distributed data.
Under edge DP constraints, we study the online change detection of community structures for CBMs. We assume an online sequence of graph observations generated from CBMs, where the underlying community structure changes at an unknown change-point. Our goal is to detect the change as quickly as possible, subject to false alarm constraints and privacy requirements [32]. Specifically, we aim to identify potential changes in the underlying community structure of CBMs based on streaming graph observations, while simultaneously ensuring the privacy of each individual graph. We propose two types of private online detection algorithms under local DP and central DP settings, respectively. Sufficient conditions for exact recovery of the unknown community structure, under DP constraints, are also established. Furthermore, we provide theoretical guarantees for the detection performance of proposed methods.
@@
The problem of private change detection is an evolving field with significant applications. It has been studied under $(\epsilon,\delta)$ differential privacy constraints, assuming parametric distributions [6, 10]. Both works assume that the pre- and post-change distributions are fully known in advance. The change detection for unknown post-change distributions was considered in [9], but the analysis was limited to univariate distributions, and thus is not directly applicable to the graph-based observations considered in our setting. In [4], online change detection for multivariate nonparametric regression was studied under local differential privacy. The proposed privacy mechanism relies on discretizing the continuous space into a set of cubes, whose number grows exponentially with the data dimension. This approach is not directly applicable to graph observations and poses computational challenges in high-dimensional settings. In [27], the authors extended the analysis of privatized networks from static to dynamic, underscoring the complexities and challenges of preserving privacy in the dynamic analysis of network data. It is worthwhile mentioning that they considered the offline change-point estimation problem with a fixed number of graph observations. In contrast, we study the online setting, where graphs arrive in a streaming fashion and the objective is to detect changes as quickly as possible. The online setting differs fundamentally from the offline one in several aspects, including its intrinsic challenges (since online methods can only use past data) and distinct performance metrics.
In the context of privacy‑preserving community detection on random graphs, Hehir et al. [22] analyzed a simple spectral approach for multi‑community settings, extending the convergence‑rate analysis of Lei & Rinaldo’s algorithm [26] and quantifying how privacy parameters influence misclassification between true and estimated labels. In our own work [35], we characterized the information‑theoretic trade‑offs of three private community‑recovery mechanisms—(a) stability‑based, (b) sampling‑based, and (c) graph‑perturbation. We showed that stability‑ and sampling‑based approaches yield a more favorable balance between the connection probabilities (i.e., random graph model parameters), accuracy and privacy budget $(\epsilon,\delta)$, albeit at higher computational cost, whereas graph‑perturbation, though cheaper, requires $\epsilon=\Omega(\log n)$ to guarantee exact recovery.
In contrast, this paper broadens the scope of analysis by considering possible network changes, in particular, shifts in community memberships. This approach presents a more comprehensive and challenging problem, as it involves not only the detection of statistical changes but also the identification of structural and compositional alterations within the network dynamics. Our methodology, therefore, addresses a broader range of changes, offering deeper insights into the complexities of dynamic network analysis.
@@
We use boldface uppercase letters to denote matrices (e.g., $\mathbf{A}$}) and boldface lowercase letters for vectors (e.g., $\mathbf{a}$}). We denote scalars by non-boldface lowercase letters (e.g., $x$), and sets by capital calligraphic letters (e.g., $\mathcal{X}$}). Let $n]\triangleq\{1,2,\cdots,n\}$ represent the set of all integers from $1$ to $n$. We use $\operatorname{Lap}(\beta)$ to denote the Laplace distribution with zero mean and scale $\beta$, and the probability density function is $\operatorname{Lap}(x|\beta)=\frac{1}{2\beta}e^{-{|x|}/{\beta}}$. For asymptotic analysis, we write the function $f(n)=o(g(n))$ when $\lim_{n\rightarrow\infty}f(n)/g(n)=0$. Also, $f(n)=\mathcal{O}(g(n))$ means there exist some constant $C>0$ such that $|f(n)/g(n)|\leq C$, $\forall n$, and $f(n)=\Omega(g(n))$ means there exists some constant $c>0$ such that $|f(n)/g(n)|\geq c,\forall n$. A summary of the key symbols is provided in Table I.
Figure 1: Overview of the proposed change‑detection scheme. We develop online community‑detection frameworks that detect changes in community memberships via CUSUM-based tests. The framework also incorporates an SDP-based community recovery component to estimate the unknown post-change communities. Both the estimation and detection components satisfy edge differential privacy to protect individual node interactions.
TABLE I: List of symbols.
$n$ Number of vertices
$\mathbf{A}$ Adjacency matrix
$\epsilon$ Privacy budget
$\delta$ Failure probability
$\bm{\sigma}^{*}$ Ground truth labels
$\hat{\bm{\sigma}}$ Estimated labels
$\mathbf{I}$ Identity matrix
$\mathbf{J}$ All ones matrix
$\mathbf{1}$ All ones vector
II Problem Setup
We begin by summarizing the problem statement in Fig. 1.
II-A Data Model
We observe a sequence of undirected graphs $\{G_{t},t\in\mathbb{N}\}$, each generated via a CBM. Each graph G=(V,E)G=({V},E) consists of $n$ vertices divided into two communities [^1] Extending our approach to encompass multiple communities is relatively straightforward, since multi-community detection can also be reformulated as an SDP problem, with analogous recovery conditions anticipated, as demonstrated in Section 5 of [21]., with community label $\bm{\sigma}^{*}=\{{\sigma}^{*}_{1},{\sigma}^{*}_{2},\cdots,{\sigma}^{*}_{n}\}$, $\sigma^{*}_{i}\in\{-1,+1\}$, $\forall i\in[n]$. The connections between vertices are represented by a weighted adjacency matrix $\mathbf{A}$, where each edge $A_{i,j}\in\{+1,-1,0\}$, $i<j$, is drawn as follows,
$\Pr(A_{i,j}=a)=\begin{cases}p(1-\zeta)&\text{if }a=\sigma^{*}_{i}\cdot\sigma^{*}_{j},\\ p\zeta&\text{if }a=-\sigma^{*}_{i}\cdot\sigma^{*}_{j},\\ 1-p&\text{if }a=0,\end{cases}$ (1)
where $p$ represents the revealed probability of an edge and $\zeta\in(0,\frac{1}{2})$. We set $A_{i,i}=0$ for all $i$ and $A_{i,j}=A_{j,i}$ for all $i>j$. This model is denoted as ${\mathsf{CBM}}(\bm{\sigma}^{*},p,\zeta)$.
Remark II.1 (Estimator of $\bm{\sigma}^{*}$).
Upon observing the adjacency matrix $\mathbf{A}$}, we may derive the maximum likelihood (ML) estimator of $\bm{\sigma}^{*}$ as (see Appendix A-B for details):
$\hat{\bm{\sigma}}_{\text{ML}}=\arg\max_{\bm{\sigma}}\{c_{1}\bm{\sigma}^{T}\mathbf{A}\bm{\sigma}+c_{0}:\sigma_{i}=\pm 1\},$ (2)
where $c_{1}=0.25\log((1-\zeta)/\zeta)$, and $c_{0}\geq 0$ is a deterministic scalar that does not depend on the labeling. It is worth mentioning that (2) is equivalent to solving the max-cut problem in the graph, which is known to be NP-hard [21]. Subsequently, we will adopt the commonly used semidefinite relaxation of the ML estimator which can be solved in polynomial time and with theoretical guarantees [21]. Define $\mathbf{Y}={\bm{\sigma}}{\bm{\sigma}}^{T}$, where $Y_{i,i}=1,\forall i\in[n]$, then the relaxed ML estimator is
$\hat{\mathbf{Y}}=\arg\max_{\mathbf{Y}}$ $\hskip 7.22743pt\text{tr}({\mathbf{A}}\mathbf{Y})\qquad\text{s.t.}\hskip 7.22743pt\mathbf{Y}\succcurlyeq\mathbf{0},\,Y_{i,i}=1,\forall i\in[n].$ (3)
We now consider a sequence of graphs $\{G_{t},t\in\mathbb{N}\}$ with $G_{t}\sim{\mathsf{CBM}}(\bm{\sigma}_{t}^{*},p,\zeta)$, and let $\mathbf{A}_{t}$ be the adjacency matrix of $G_{t}$. We assume there is an abrupt change to the community label at time $\nu$, which is called a change-point and is unknown. The problem can be formulated as follows,
[…]
where $\bm{\sigma}^{\text{pre}}$ and $\bm{\sigma}^{\text{post}}$ represent the pre- and post-change communities, respectively. We further assume $\bm{\sigma}^{\text{pre}}$ is known since it can usually be estimated from historical data under the pre-change regime. The post-change community $\bm{\sigma}^{\text{post}}$ is assumed to be unknown. We measure the strength of change using the Hamming distance between $\bm{\sigma}^{\text{pre}}$ and $\bm{\sigma}^{\text{post}}$,
$\operatorname{Ham}({\bm{\sigma}^{{\text{pre}}}},{\bm{\sigma}^{{\text{post}}}}):=\sum\nolimits_{i=1}^{n}\mathbb{I}\{\sigma_{i}^{{\text{pre}}}\neq\sigma_{i}^{{\text{post}}}\}.$
We note that $\operatorname{Ham}({\bm{\sigma}^{{\text{pre}}}},{\bm{\sigma}^{{\text{post}}}})\leq n/2$, which holds without loss of generality since larger distances can be reduced by flipping community labels.
Our goal is to design a detection algorithm that is able to protect the user’s privacy while detecting the change as quickly as possible from streaming data. The detection is performed through a stopping time $\tau$ on the observation sequence at which the change is declared [32]. To measure the performance of a stopping time, we introduce two metrics commonly used in online detection problems: detection delay and false alarm performances. Let $\Pr_{\nu}$ denote the probability measure on the observation sequence when the change-point is $\nu$, and let $\mathbb{E}_{\nu}$ denote the corresponding expectation. Also, $\Pr_{\infty}$ and $\mathbb{E}_{\infty}$ denote the probability and expectation operator when there is no change (i.e., $\nu=\infty$). The commonly adopted definition of worst-case average detection delay (WADD) is [29]:
where $\{\mathcal{F}_{k},\ k\in\mathbb{N}\}$ is the filtration $\mathcal{F}_{k}=\sigma(G_{1},\ldots,G_{k})$, with $\mathcal{F}_{0}$ denoting the trivial sigma algebra. We measure the false alarm performance of a detector (stopping time) $\tau$ in terms of its Average Run Length (ARL), defined as $\mathbb{E}_{\infty}\left[\tau\right]$. The larger the ARL, the smaller the probability of false alarms. We aim to design a detection method that satisfies the desired requirement on ARL while having a small WADD, under the private setting to be detailed next.
II-B Differential Privacy for Graphs
In addition to the dynamic community detection, we require that the algorithm also protect the individual’s data privacy. In the context of graph data, our threat model is structured around the principle of edge differential privacy (DP), incorporating both local and central settings.
Local DP (LDP) offers robust privacy guarantees during the data collection process, ensuring that individual data contributions are privatized at the source. This is crucial in scenarios where trust in the data curator is limited. Similar to [27], we consider the privacy mechanism which privatizes each edge $A_{i,j}$ to $\tilde{A}_{i,j}$ following a conditional distribution $Q$ as defined below.